# On the inverse Möbius transformation and unrestricted partitions

### 10 Feb 2024 v1

@@ 1. Introduction
Let $a_{n}$ and arbitrary sequence of real or imaginary terms and $n$ a positive integer number, Then
(2.4) $\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}a_{l(\pi)-s(\pi)+j}=\sum_{d|n}(-1)^{d}(-a_{1}+a_{2}-a_{3}+\cdots+(-1)^{d}a_{d})$
@@
Let $a_{n}$ and arbitrary sequence of real or imaginary terms and $n$ a positive integer number, Then
(2.5) $\begin{split}\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}(-1)^{l(\pi)-s(\pi)+j}a_{l(\pi)-s(\pi)+j}\\ &=\sum_{d|n}(-1)^{d}(a_{1}+a_{2}+a_{3}+\cdots+a_{d})\end{split}$
@@
Let $a_{n}$ and arbitrary sequence of real or imaginary terms with $a_{0}=0$.
(2.6) $\begin{split}\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)}&\sum_{j=1}^{s(\pi)}\frac{a_{l(\pi)-s(\pi)+j}+a_{l(\pi)-s(\pi)+j-1}}{2}=\sum_{d|n}a_{d}\end{split}$
We should consider that $a_{0}=0$ can be defined for any sequence $a_{1},a_{2},\cdots$, simply by appending zero to it: $0,a_{1},a_{2},\cdots$.
@@
If in Theorem 2.4 $a_{n}=n^{2}$, Then
(2.7) $\begin{split}\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}(l(\pi)-s(\pi)+j)^{2}\\ &=\sum_{d|n}\frac{d(d+1)}{2}\\ &=\frac{\sigma_{2}(n)+\sigma_{1}(n)}{2}\end{split}$
@@
If in Theorem 2.4 $a_{n}=n$, Then
(2.8) $\begin{split}\sum_{\pi\in P(n)}t^{k(\pi)-1}2^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}l(\pi)-s(\pi)+j\\ &=\frac{\sigma_{1}(n)+\tau_{odd}(n)}{2}\end{split}$
@@
If in the Corollary 2 $a_{2k}=0,a_{2k+1}=1$ (for $k\geq 0$), Then
(2.9) $\begin{split}\sum_{\pi\in P(n)}t^{k(\pi)-1}2^{Q(\pi)-1}s(\pi)=\tau_{odd}(n)\\ \end{split}$
@@ 3. Proof of the Theorems and Corollaries
If we set $t=-1$ , $u=2$, and we use the Heine’s transformation as Andrews in [26],
@@
$(-2)\sum_{n\geq 1}a_{n}$ $q^{n}\frac{\left(q\right)_{n-1}}{(-q)_{n}}\sum_{i\geq 0}\frac{(-q)_{i}\left(q^{n}\right)_{i}}{(-q^{n+1})_{i}\left(q\right)_{i}}q^{i}=$
$(-2)\sum_{n\geq 1}a_{n}q^{n}\frac{\left(q\right)_{n-1}}{(-q)_{n}}\frac{(-q)_{\infty}(q^{n+1})_{\infty}}{(-q^{n+1})_{\infty}(q)_{\infty}}\sum_{j\geq 0}\frac{(q)_{j}\left(q^{n}\right)_{j}}{(q)_{j}\left(q^{n+1}\right)_{j}}(-q)^{j}$
$(-2)\sum_{n\geq 1}a_{n}q^{n}\frac{\left(q\right)_{n-1}}{(-q)_{n}}\frac{(-q)_{n}}{(q)_{n}}\sum_{j\geq 0}\frac{(1-q^{n})}{(1-q^{n+j})}(-q)^{j}$
$(-2)\sum_{n\geq 1}a_{n}\sum_{j\geq 0}\frac{(-1)^{j}(q)^{n+j}}{(1-q^{n+j})}$
$=(-2)a_{1}\left[\frac{(-1)^{0}q}{1-q}+\frac{(-1)^{1}q^{2}}{1-q^{2}}+\frac{(-1)^{2}q^{3}}{1-q^{3}}\cdots\right]+$
$(-2)a_{2}\left[\frac{(-1)^{0}q^{2}}{1-q^{2}}+\frac{(-1)^{1}q^{3}}{1-q^{3}}+\frac{(-1)^{2}q^{4}}{1-q^{4}}\cdots\right]+$
$(-2)a_{3}\left[\frac{(-1)^{0}q^{3}}{1-q^{3}}+\frac{(-1)^{1}q^{4}}{1-q^{4}}+\frac{(-1)^{2}q^{5}}{1-q^{5}}\cdots\right]+$
@@
$=(-2)\sum_{k=1}\left(\sum_{i\geq 1}^{k}(-1)^{k-i}a_{i}\right)\frac{q^{k}}{1-q^{k}}$
$=(-2)\sum_{k=1}q^{k}\sum_{d|k}\left(\sum_{i\geq 1}^{d}(-1)^{d-i}a_{i}\right)$
$=(-2)\sum_{k=1}q^{k}\sum_{d|k}(-1)^{d}\sum_{i\geq 1}^{d}(-1)^{i}a_{i}$
$=(-2)\sum_{k=1}q^{k}\sum_{d|k}(-1)^{d}(-a_{1}+a_{2}-a_{3}+\cdots+(-1)^{d}a_{d})$

### 22 Feb 2024 v2

@@ 1. Introduction
Let $a_{n}$ and arbitrary sequence of real or imaginary terms , $t$ a real or imaginary number different of $1$ or 00, and $n$ a positive integer number, Then
(2.4) $\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}a_{l(\pi)-s(\pi)+j}=\sum_{d|n}(t)^{d}\left(\frac{a_{1}}{t}+\frac{a_{2}}{t^{2}}+\frac{a_{3}}{t^{3}}+\cdots+\frac{a_{d}}{t^{d}}\right)$
@@
Let $a_{n}$ and arbitrary sequence of real or imaginary terms , $t$ a real or imaginary number different of $1$ or 00, and $n$ a positive integer number, Then
(2.5) $\begin{split}\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}(t)^{l(\pi)-s(\pi)+j}a_{l(\pi)-s(\pi)+j}\\ &=\sum_{d|n}(t)^{d}(a_{1}+a_{2}+a_{3}+\cdots+a_{d})\end{split}$
@@
Let $a_{n}$ and arbitrary sequence of real or imaginary terms with $a_{0}=0$ and $t$ a real or imaginary number different of $1$ or 00,.
(2.6) $\begin{split}\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}\left(a_{l(\pi)-s(\pi)+j}-ta_{l(\pi)-s(\pi)+j-1}\right)=\sum_{d|n}a_{d}\end{split}$
We should consider that $a_{0}=0$ can be defined for any sequence $a_{1},a_{2},\cdots$, simply by appending zero to it: $0,a_{1},a_{2},\cdots$. This Corollary show that the Mobius inverse transformation of any sequence $a_{n}$ with $a_{0}=0$ could be interpreted as weighted partition. We should notice that the left therm depend in $t$ and the right term not.
@@
If in Theorem 2.4 $a_{n}=n$ and $t=-1$, Then
(2.7) $\begin{split}\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}l(\pi)-s(\pi)+j\\ &=\sum_{d|n}(-1)^{d}\left(\frac{1}{-1}+\frac{2}{(-1)^{2}}+\frac{3}{(-1)^{3}}+\cdots+\frac{d}{(-1)^{d}}\right)\\ &=\sum_{d|n}(-1)^{d}\left(-1+2-3+\cdots+(-1)^{d}d\right)\\ &=\sum_{d|n}\left[\frac{d}{2}+\left(\frac{1-(-1)^{d}}{2}\right)\frac{1}{2}\right]\\ &=\frac{\sigma_{1}(n)+\tau_{odd}(n)}{2}\end{split}$
@@
If in Theorem 2.4 $a_{n}=n^{2}$ and $t=-1$, Then
(2.8) $\begin{split}\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}&\sum_{j=1}^{s(\pi)}(l(\pi)-s(\pi)+j)^{2}\\ &=\sum_{d|n}(-1)^{d}\left(\frac{1^{2}}{-1}+\frac{2^{2}}{(-1)^{2}}+\frac{3^{2}}{(-1)^{3}}+\cdots+\frac{d^{2}}{(-1)^{d}}\right)\\ &=\sum_{d|n}(-1)^{d}\left(-1+4-9+\cdots+(-1)^{d}d^{2}\right)\\ &=\sum_{d|n}\frac{d(d+1)}{2}\\ &=\frac{\sigma_{2}(n)+\sigma_{1}(n)}{2}\end{split}$
@@
If in the Corollary 2 $a_{2k}=0,a_{2k+1}=1$ (for $k\geq 0$) and $t=-1$, Then
(2.9) $\begin{split}&\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}1=\sum_{d|n}\frac{1-(-1)^{d}}{2}\\ &\sum_{\pi\in P(n)}(-1)^{k(\pi)-1}2^{Q(\pi)-1}s(\pi)=\tau_{odd}(n)\\ \end{split}$
@@ 3. Proof of the Theorems and Corollaries
If we set $t=t$ , $u=(\frac{t-1}{t})$, and we use the Heine’s transformation as Andrews in [26],
@@
$(tu)\sum_{n\geq 1}a_{n}$ $q^{n}\frac{\left(q\right)_{n-1}}{(tq)_{n}}\sum_{i\geq 0}\frac{(tq)_{i}\left(q^{n}\right)_{i}}{(tq^{n+1})_{i}\left(q\right)_{i}}q^{i}=$
$(tu)\sum_{n\geq 1}a_{n}q^{n}\frac{\left(q\right)_{n-1}}{(tq)_{n}}\frac{(tq)_{\infty}(q^{n+1})_{\infty}}{(tq^{n+1})_{\infty}(q)_{\infty}}\sum_{j\geq 0}\frac{(q)_{j}\left(q^{n}\right)_{j}}{(q)_{j}\left(q^{n+1}\right)_{j}}(tq)^{j}$
$(tu)\sum_{n\geq 1}a_{n}q^{n}\frac{\left(q\right)_{n-1}}{(tq)_{n}}\frac{(tq)_{n}}{(q)_{n}}\sum_{j\geq 0}\frac{(1-q^{n})}{(1-q^{n+j})}(tq)^{j}$
$(tu)\sum_{n\geq 1}a_{n}\sum_{j\geq 0}\frac{(t)^{j}(q)^{n+j}}{(1-q^{n+j})}$
$=(tu)a_{1}\left[\frac{(t)^{0}q}{1-q}+\frac{(t)^{1}q^{2}}{1-q^{2}}+\frac{(t)^{2}q^{3}}{1-q^{3}}\cdots\right]+$
$(tu)a_{2}\left[\frac{(t)^{0}q^{2}}{1-q^{2}}+\frac{(t)^{1}q^{3}}{1-q^{3}}+\frac{(t)^{2}q^{4}}{1-q^{4}}\cdots\right]+$
$(tu)a_{3}\left[\frac{(t)^{0}q^{3}}{1-q^{3}}+\frac{(t)^{1}q^{4}}{1-q^{4}}+\frac{(t)^{2}q^{5}}{1-q^{5}}\cdots\right]+$
@@
$=(tu)\sum_{k=1}\left(\sum_{i\geq 1}^{k}(t)^{k-i}a_{i}\right)\frac{q^{k}}{1-q^{k}}$
$=(tu)\sum_{k=1}q^{k}\sum_{d|k}\left(\sum_{i\geq 1}^{d}(t)^{d-i}a_{i}\right)$
$=(tu)\sum_{k=1}q^{k}\sum_{d|k}(t)^{d}\sum_{i\geq 1}^{d}(t)^{-i}a_{i}$
$=(tu)\sum_{k=1}q^{k}\sum_{d|k}(t)^{d}\left(\frac{a_{1}}{t}+\frac{a_{2}}{t^{2}}+\frac{a_{3}}{t^{3}}+\cdots+\frac{a_{d}}{t^{d}}\right)$
which conclude the proof. ∎
Proof of Corollary 1
Proof.
In effect, If in the Theorem 2.4 we consider an auxiliary sequence $b_{n}$ defined by
(3.6) $a_{n}=(t)^{n}b_{n}$
replacing $a_{n}$ in the Theorem 2.4 we got
(3.7) $\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}(t)^{l(\pi)-s(\pi)+j}b_{l(\pi)-s(\pi)+j}=\sum_{d|n}(t)^{d}(b_{1}+b_{2}+b_{3}+\cdots+b_{d})$
which conclude the proof. ∎
Proof of Corollary 2
Proof.
If we define the auxiliary sequence $b_{n}$ defined as
(3.8) $b_{n}=a_{1}+a_{2}+\cdots+a_{n}$
with $b_{0}=0$, then
(3.9) $a_{n}=b_{n}-b_{n-1}$
for $n\geq 1$
In the Corollary 1, we conclude then,
(3.10) $\begin{split}&\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}(t)^{l(\pi)-s(\pi)+j}\left[b_{l(\pi)-s(\pi)+j}-b_{l(\pi)-s(\pi)+j-1}\right]\\ &=\sum_{d|n}(t)^{d}b_{d}\end{split}$
If we then define an auxiliary sequence $c_{n}$ as
$b_{n}=(t)^{-n}c_{n}$
Then, the interior sum is
(3.11) $\begin{split}&\sum_{j=1}^{s(\pi)}(t)^{l(\pi)-s(\pi)+j}\left[(t)^{-(l(\pi)-s(\pi)+j)}c_{l(\pi)-s(\pi)+j}-(t)^{-(l(\pi)-s(\pi)+j-1)}c_{l(\pi)-s(\pi)+j-1}\right]\\ &\sum_{j=1}^{s(\pi)}\left[c_{l(\pi)-s(\pi)+j}-tc_{l(\pi)-s(\pi)+j-1}\right]\\ \end{split}$
Finally,
(3.12) $\begin{split}&\sum_{\pi\in P(n)}(t)^{k(\pi)-1}\left(\frac{t-1}{t}\right)^{Q(\pi)-1}\sum_{j=1}^{s(\pi)}\left[c_{l(\pi)-s(\pi)+j}-tc_{l(\pi)-s(\pi)+j-1}\right]\\ &=\sum_{d|n}c_{d}\end{split}$